A train is travelling at 80km/h. how far does the train go in 1 minute?
Answer by fcabanski(1391)

There are 60 minutes per hour. So divide by 60 to find km/minute.


80 km / h*60m/h = 80/60 km/m = 1.33 km/m.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)